OpenStudy (anonymous):

Is there anyone in FLVS that can help me with Algebra 1 3.04?

OpenStudy (anonymous):

3. A textbook company keeps track of its daily cost for printing whole textbooks to different schools, but can only print up to 500 books a day. The daily cost (C) to print x number of textbooks is given by C(x) = 3.50x + 1200. 1. Using the function, determine one possible daily cost by evaluating the function for any input of your choice. 3.5 x (50) + 1200 175 + 1200 1375 2. For what value of x does C(x) = 1900? 362.86 3. Interpret the statement C(20) = 1270. It cost $1270 to print 20 books. 4. Interpret the statement C(w) = p. 5. Explain whether C(40.5) = 1341.75 is correct and possible given the context of this problem. 3. C(x) is the cost for printing x books. This line means that "It cost's $1270 to print 20 books" 4. Same thing as 3: It costs p dollars to print w books 5: Can you print half a book? No, so C(40.5) doesn't make sense in the context. 4. Bill is able to save $35/week after working part-time and paying his expenses. These two formulas show his weekly savings: 1. f(1) = 35, f(n) = f(1) + f(n-1) for n > 1 2. f(n) = 35n 1. Which one of these formulas show the sequence written recursively, and which shows it written explicitly? Justify your explanations. 2. Use the recursive formula to make a table of values for 1 ≤ n ≤ 5. Show your calculations. Explain what your table means. 3. Use any formula of your choice to find f(40). Explain why you chose that method and what your answer means. Show your calculations. 4. Given the sequence of numbers: 5, 6, 8, 11, 15, 20, 26, 33, 41,… Explain whether or not this sequence can be considered a function.   5. Create your own equation written in standard form. 1. Determine any point that is a solution and will be on this line. Justify your answer algebraically. 2. Explain how to find the x- and y- intercepts using complete sentences. Write the intercepts as ordered pairs. 3. Rewrite the equation using function notation. Graph the function and show the intercepts.   6. Compare the graph from question 2 and the equation from question 5. How are the graphs similar and different? Determine which has the greater y-value when x is 160, and explain how you determined it.

rishavraj (rishavraj):

2. For what value of x does C(x) = 1900? 362.86 its not tht C(x) = 3.5x + 1200 = 1900 get x

OpenStudy (anonymous):

342.857143?? am I supposed to divide 1200 by 3.5

rishavraj (rishavraj):

see 3.5x + 1200 = 1900 3.5x = 700 x =??

OpenStudy (anonymous):

ohh 200

rishavraj (rishavraj):

yup